OpenStudy (anonymous):

10xsquared+6x-28 expression

OpenStudy (anonymous):

10x^2+6x-28 =2(5x^2+3x-14) =2(5x-7)(x+2) 5x-7=0 --->x=7/5 x+2=0 --->x=-2

OpenStudy (anonymous):

the answers are a (5x+7)(2x-4) b 2(5x-7)(x+4) c 2(5x-7)(x+2) d (5x-7)(2x+4)

OpenStudy (anonymous):

answer would be C) 2(5x-7)(x+2)

OpenStudy (anonymous):

are you good at this stuff

OpenStudy (anonymous):

Yea, a little. I just finished up Differential Equation for my Electrical Engineering major. This stuff comes easy after a while and pracitve

OpenStudy (anonymous):

wel i go to the institute of photography

OpenStudy (anonymous):

and i have to place out of math so far not looking good

OpenStudy (anonymous):

Lol..probably not quite as mathematically demanding as this stuff, but that makes you the lucky one I'd say!

OpenStudy (anonymous):

errrrr

OpenStudy (anonymous):

Get a tutor, they don't cost much for algebra and things like that. Then again, I'm in Texas, everything is cheaper here.

OpenStudy (anonymous):

i dont need it to take great pics but my degree does lol

OpenStudy (anonymous):

im 35 its been a while lol

OpenStudy (anonymous):

lol..I didn't think you were that old from your picture. I thought you were in college when you said institute of photography

OpenStudy (anonymous):

I meant college "age"...20's

OpenStudy (anonymous):

i am in collage but thanks good skin i guess lol

OpenStudy (anonymous):

haha..it helps..anyways, i'm gonna get going..you need anymore help?

OpenStudy (anonymous):

taking the placement tommorow

OpenStudy (anonymous):

tomorrow

OpenStudy (anonymous):

yikes, well good luck on that, I hope my two problems come through and help you out on your test tomorrow.

OpenStudy (anonymous):

thanks the crappy thing is if i score high i dont even have to take the math

OpenStudy (anonymous):

Not taking math is a good thing..I've done alebra 1 and 2, calculus 1, 2, and 3 and just finished differential equations and still have 2 more classes.

OpenStudy (anonymous):

wow when i was in school i had algebra 1 2 geometry i dont remember a thing lol

OpenStudy (anonymous):

i just want to take awsome pictures forget the math but thanks

OpenStudy (anonymous):

good luck with whatever you do ;)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ive got one i cant figure how to write

OpenStudy (anonymous):

lol, what is it?

OpenStudy (anonymous):

y wierd symbol z-little4over ylittle-2zlittle2 lolol

OpenStudy (anonymous):

lol..i can't even decipher what it might look like...

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\]

OpenStudy (anonymous):

is it that kind of weird symbol?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol..there are over 300 weird symbols in math.

OpenStudy (anonymous):

y looks like a little z

OpenStudy (anonymous):

\[\le \ge ><\neq \approx\]

OpenStudy (anonymous):

a cursive little z?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

\[\zeta \xi\]

OpenStudy (anonymous):

you have e mail

OpenStudy (anonymous):

yes i do..haha..this is amusing. jberry08@tamu.edu

OpenStudy (anonymous):

right lol i sent it

OpenStudy (anonymous):

waiting now..email server is slow or something give me a few minutes.

OpenStudy (anonymous):

i have an edu too cristinashields @stu.aii.edu

OpenStudy (anonymous):

lol..yea, .edu is used for "educational" things like .com is used for "commercial" things

OpenStudy (anonymous):

you put jberry"zero-eight" right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

weird it still hasn't come through

OpenStudy (anonymous):

perhaps try my google account.. jberry08@gmail.com

OpenStudy (anonymous):

cool im doing it from my phone lol

OpenStudy (anonymous):

whatever works, i do alot of emailing on my phone too..makes me feel businessman like

OpenStudy (anonymous):

galexy mesmerize i took a pic of it lol

OpenStudy (anonymous):

what is galaxy mesmerize?

OpenStudy (anonymous):

phone

OpenStudy (anonymous):

oh, aren't you cool..

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

yea same here unfortunately

OpenStudy (anonymous):

jp

OpenStudy (anonymous):

we're pretty cool..we're both on a tutoring math site waiting for an email

OpenStudy (anonymous):

you get it

OpenStudy (anonymous):

no nothing has come through..there must be a problem..i would say text me the picture but that's weird and then i'd have to change my number

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

im not a stalker

OpenStudy (anonymous):

lol..well if you want to 4325539308

OpenStudy (anonymous):

you get that one

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

one second let me look at it

OpenStudy (anonymous):

did it make sense now that you see it lol

OpenStudy (anonymous):

\[(y ^{3}z ^{-4})/(y ^{-2}z ^{2})\]

OpenStudy (anonymous):

is that what it said??

OpenStudy (anonymous):

idk if that was a 3 it looked like a t with a little o at the bottom . i dk

OpenStudy (anonymous):

or a wierd a

OpenStudy (anonymous):

yea i think it is a three.. none of the answer are correct though. It should be: \[y ^{5}/z ^{6}\] the first answer is closest to that..but it is multiplied by itself again..idk..this question is dumb and incorrect

OpenStudy (anonymous):

i took a picture of it right off my test i have a chance to retake

OpenStudy (anonymous):

well i'm good at simplifying exponents and so is my calculator and the right answer is not given.. if i had to choose i'd go with A

OpenStudy (anonymous):

I sent you the solution to both problems..hope they help, i know that the first problem is all screwed up though..the answer isn't there..must be a mistake

OpenStudy (anonymous):

its messed up if the answer isint there lol

OpenStudy (anonymous):

yea..kind of rediculous for a test to have incorrect answers..>:/

OpenStudy (anonymous):

did you get those texted pics??

OpenStudy (anonymous):

how about i pay u you take it for me lol

OpenStudy (anonymous):

lol..idk if i could get to wherever that test is by tomorrow..

OpenStudy (anonymous):

and i'm kinda missing a head full of blonde hair

OpenStudy (anonymous):

i have a week

OpenStudy (anonymous):

..i thought you were taking it tomorrow??

OpenStudy (anonymous):

its online lol

OpenStudy (anonymous):

haha..that would be quite dishonest..how much??

OpenStudy (anonymous):

ive got 3 to take

OpenStudy (anonymous):

i took it and got a 75 booo

OpenStudy (anonymous):

idk if i could do that..i won't be near a computer next week..i could do it tonight..that's about all i can do, i'm busy after tonight..

OpenStudy (anonymous):

so i have a redo

OpenStudy (anonymous):

well i get a b or above 100$

OpenStudy (anonymous):

haha..okay. how do they know if it's you or not?

OpenStudy (anonymous):

they dont you log in my campuss acct and take it

OpenStudy (anonymous):

you good at excel

OpenStudy (anonymous):

can it be done tonight? and yea i'm good with excel

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

okay..so how will i get my $100??

OpenStudy (anonymous):

you tell me where to post it when they give my grade to me